37. To determine the heat capacity of a solution calorimeter (constant pressure calorimeter), 75.3g of copper at 370.5 K was placed into 100.3g of water in the calorimeter. The initial temperature of the water was 294.4 K; the final temperature of the system was 298.8 K. What is the heat capacity of the calorimeter? For Cu, specific heat -0.385 J/Kg; for water, specific heat- 4.184 J/Kg. 38. The same calorimeter used in #37 was used to find the heat flow associated with the dissolution of 24g of NaCl in 176g of water. The observed ΔΤ (for both the calorimeter and the solution) was -1.6K. The calorimeter heat capacity is 60 J/K. The specific heat of the resulting NaCl solution is 3.64 J/Kg. What is the heat flow for this process?  

38. The same calorimeter used in #37 was used to find the heat flow associated with the dissolution of 24g of NaCl in 176g of water. The observed ∆T (for both the calorimeter and the solution) was -1.6K. The calorimeter heat capacity is 60 J/K. The specific heat of the resulting NaCl solution is 3.64 J/Kg. What is the heat flow for this process?
Answer #1

38) q = Ccal * ΔT + m * cp * ΔT

We calculate the mass involved:

m = 24 g NaCl + 176 g H2O = 200 g

cp = 3.64 J / g * K

ΔT = -1.6 K

Ccal = 60 J / K

We substitute and calculate:

q = 60 * (- 1.6) + 200 * 3.64 * (-1.6) = -1260.8 J.